MASTERKIY:

Diego’s family car holds 14 gallons of fuel. Each day the car uses 0.6 gallons of fuel. A warning light comes on when the remaining fuel is 1.5 gallons or less. d. Write an equation to determine the number of days, t, Diego’s father can drive the car without the warning light coming on. A. Type the equation in the space below. B. Solve the equation and round your answer to the nearest whole number.

Extrinix:

Ok, so first off, remove the amount of gas that the warning light activates, 14 - 1.5 = 12.5 gallons Now, if the car uses 0.6 gallons per day, you would subtract that from the total amount: \(12.5 \times 0.6 = t\) < so this would be your equation, now all you do is solve it, \(12.5 \times 0.6 = ? \rightarrow 7.5\) days.

darkknight:

extrinix wrote:
Ok, so first off, remove the amount of gas that the warning light activates, 14 - 1.5 = 12.5 gallons Now, if the car uses 0.6 gallons per day, you would subtract that from the total amount: \(12.5 \times 0.6 = t\) < so this would be your equation, now all you do is solve it, \(12.5 \times 0.6 = ? \rightarrow 7.5\) days.
"Ok, so first off, remove the amount of gas that the warning light activates, 14 - 1.5 = 12.5 gallons" this part i agree with "Now, if the car uses 0.6 gallons per day, you would subtract that from the total amount:" not sure what you meant here but moving on 12.5 times .6 is not what you do Unit analysis, 12.5 is gallons, .6 is a rate of gallons per day. Sometimes its hard to know what you need to do but if you multiply 12.5 and 6 you are multiplying gallons with gallons per day which results in gallons^2 per day. You want to find the number of days not gallons squared per day